OpenStudy (anonymous):

Write the expression below as a single logarithm: (1/3)Inx + 4In(yz) - 3In(x^2yz). How do I begin? Write the expression below as a single logarithm: (1/3)Inx + 4In(yz) - 3In(x^2yz). How do I begin? @Mathematics

OpenStudy (anonymous):

you move the numbers in front to become powers of the expressions you're taking ln of. i.e. \[blna=\ln(a^b)\]

OpenStudy (anonymous):

(1/3)^x + 4^yz - 3^x^2yz Is that what you mean?

jhonyy9 (jhonyy9):

yes is right but you know again that lna +lnb=ln(a*b)

OpenStudy (anonymous):

no, \[(1/3) \ln x = \ln [x^(1/3)]\] and so on. its a logarithm law :)

jhonyy9 (jhonyy9):

and lna - lnb=ln(a/b)

OpenStudy (anonymous):

oops, equation messed up lol. correction: \[{1/3}\ln x =\ln [x^{(1/3)}]\]

OpenStudy (anonymous):

you do that for all three logs, and then use the laws that jhonyy9 mentioned to rewrite the expression as a single logarithm to the base e :)

OpenStudy (anonymous):

\[ In[x^(1/3)] + In[(yz)^4] - In[(x^2yz)^3]\]

OpenStudy (anonymous):

yup thats right :)

jhonyy9 (jhonyy9):

ln((cuberootx)(yz)4)/(x2yz)3) = ln(cuberoot(x)(yz)/(x2))

OpenStudy (anonymous):

now use the fact that ln x +ln y = ln xy and ln x - ln y = ln (x/y)

OpenStudy (anonymous):

\[In[x^(1/3)] + In[(yz)^4] = In[(x^2yx)^3]\] There's the first one.

OpenStudy (anonymous):

unfortunately, you can't set the one side equal the other, because the whole expressions isn't equal to zero.

OpenStudy (anonymous):

I have never struggled so much in a math class before.

OpenStudy (anonymous):

what class is it?

OpenStudy (anonymous):

It's called College Algebra, but is actually Pre-Calculus.

OpenStudy (anonymous):

oh that class was tough ><

OpenStudy (anonymous):

Tell me about it.

OpenStudy (anonymous):

it gets better though, once you get the hang of everything :)

OpenStudy (anonymous):

I thought that I was going to happen, but now I'm in my final week

OpenStudy (anonymous):

ouch. well on the bright side, you won't have to deal with this anymore after the week is through :D

OpenStudy (anonymous):

lol. Yeah, with the possibility of having to take this over again! I'm really looking forward to it. :D

OpenStudy (anonymous):

i wish you luck with that possibility :)